# Pure Moving Electromagnetic Media Consisting of Magnetochiral Metasurfaces

### v2

@@
Abstract
@@
Numerical studies demonstrate that magneto-chiral (MCh) metasurfaces consisting of double Z-type gammadions with perpendicularly magnetized substrate exhibit bianisotropy for microwaves. Effective polarizability tensors of the metasurfaces, extracted from the reflection and transmission coefficients, have components corresponding to non-reciprocal moving-type bianisotropy as well as reciprocal chiral-type bianisotropy and non-reciprocal magneto-optical effects. The combination of the metasurfaces and contraposition MCh metasurfaces cancels the chiral-type bianisotropy and magneto-optical effect, leading to pure moving electromagnetic media. Achieved perfect transmission with phase difference of $\pi$ realizes an ideal gyrator for arbitrary spatial and polarization modes.
preprint: APS/123-QED
I Introduction
@@
Symmetries relate to physical laws. Noether’s theorem says that a system with a continuous symmetry has a conserved quantity, for example, the momentum conservation for spatially uniform systems and the energy conservation for the time-independent systems. Another kind of symmetries is a discrete one, such as spatial inversion and time-reversal symmetry. These symmetries and laws impose restrictions on physical phenomena, and play an essential role in basic and applied physics. Therefore, symmetry breaking would result in emergence of a vast variety of phenomena [1]. In electromagnetism, a breaking of space-inversion and/or time-reversal symmetry allows bianisotropy [2, 3], which represents cross-correlated electromagnetic inductions among electric polarizability, magnetization, and electromagnetic fields in constitutive equations, i.e., magnetoelectric coupling. The bianisotropy is classified into the four categories associated with reciprocity and polarization dependence as follows [4]: a reciprocal chiral-type with polarization plane rotation, a reciprocal omega-type without polarization rotation, a non-reciprocal Tellegen-type with polarization rotation, and a non-reciprocal moving-type without polarization rotation.
While the control of the bianisotropy is a challenge in natural materials with broken symmetry such as multiferroic materials [5, 6, 7], man-made structured materials, referred to as metamaterials, are suitable for investigation and control in application of bianisotropy [8, 9, 10, 11, 12]. In particular, metasurfaces with two-dimensional planar structures are good playgrounds to embody complex bianisotropic media [13, 14, 15, 16]. Here we focus on magnetochiral (MCh) metasurfaces to realize the fourth type of the bianisotropy classified above, namely the non-reciprocal moving type. The MCh metasurfaces are composed of both chiral meta-atoms with broken space-inversion symmetry and magnetic meta-atoms with broken time-reversal symmetry. The optical MCh effect is manifested in a transmission coefficient difference, which is dependent on propagation direction but independent of polarization, referred to as directional birefringence [17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33]. The directional birefringence corresponds to the moving-type bianisotropy.
@@
In this work, we numerically calculate transmission and reflection coefficient spectra at microwave frequencies to evaluate effective polarizability tensor [34, 35, 36] of MCh metasurfaces. The effective polarizability tensor, extracted from the transmission and reflection coefficients at normal incidence, clearly show optical activity, magneto-optical effect, and electromagnetic moving-medium effect. Furthermore, a composite metasurface consisting of the ordinary and contraposition MCh metasurcfaces realizes pure moving electromagnetic media without optical activity and magneto-optical effect. The pure moving media demonstrate the perfect transmission with a phase difference of $\pi$ depending on propagation direction, bringing about an ideal gyrator [37].
@@
The pure moving electromagnetic media are of great interest in terms of development of functional metasurfaces. A similar story can be found in the development of the well-known negative refractive index metamaterials consisting of double split-ring resonators (SRRs). A single C-shaped SRR shows the omega-type bianisotropy due to electromagnetic interaction in addition to magnetic permeability [38, 39]. This brings about electromagnetic responses not only to magnetic field but also to electric field. An inverse-C-shaped SRR is thus incorporated in the inner circle of the C-shaped SRR, resulting in vanishment of the omega-type electromagnetic interaction but remaining permeability component. In this way, the double SRRs construct pure negative magnetic permeability media, leading to negative refractive index metamaterials. Similarly, the pure moving medium obtained in this study is a key to realize an ideal gyrator for arbitrary spatial and polarization modes. Applications of the ideal gyrator include an isolater, circulator, non-reciprocal holography [40], two-wavelength lasing [28], and artificial gauge field for light [41]. The pure moving electromagnetic medium is also relevant in basic physics in an analogy with a hypothetical particle called dyon [42], and from the view point of interplay between physics and algebraic geometry in Fresnel-Kummer surfaces [43, 44].
@@
This paper is organized in eight sections. Section II describes the theoretical background of effective polarizability tensor, and explains bianisotropy in four categories. Section III details numerical calculation setup. Section IV devotes to calculation results of microwave transmission and reflection coefficient spectra of uniaxial bianisotropic MCh metasurfaces consisting of double Z-type gammadion [45] with perpendicularly magnetized substrate. The extracted effective polarizability tensors of chiral and MCh metasurfaces are illustrated in Sections V and VI, respectively. Section VII presents pure moving electromagnetic medium by introducing contraposition MCh metasurfaces with the inverse chirality and converse magnetization of the ordinary MCh metasurfaces. The pure moving medium brings about perfect transmission and phase difference of $\pi$. Section VIII concludes the paper.
II Theoretical background
II.1 Effective polarizability tensor
@@
The electromagnetic properties of a metasurface, which interacts with normally incident plane waves propagating in the $+z$ direction, can be fully characterized by a following effective polarizability tensor $\hat{\alpha}_{ij}$ defined as
@@
Here we write the plane waves to be, $E,H\propto{\rm e}^{j(-\vec{k}\cdot\vec{r}+\omega t)}$, where $j=\sqrt{-1}$ is an imaginary unit. In this notation, the effective polarizability tensor components can be expressed in a non-dimensional manner to be

### v3

# Pure moving optical media consisting of magnetochiral metasurfaces

@@
Pure moving optical media consisting of magnetochiral metasurfaces
[…]
journal: opticajournalarticletype: Research Article
{abstract*}
@@
We report numerical studies of microwave bianisotropies by magnetochiral (MCh) metasurfaces consisting of double Z-type gammadions with perpendicularly magnetized substrates. The metasurfaces’ effective polarizability tensor, extracted from calculated reflection and transmission coefficients, has components of the non-reciprocal moving-type bianisotropy as well as the reciprocal chiral-type bianisotropy and non-reciprocal magneto-optical (MO) effect. Combination of the metasurfaces with contraposition MCh metasurfaces cancels both the chiral-type bianisotropy and MO effect, resulting in pure moving optical media. Achieved perfect transmission with phase difference of $\pi$ realizes an ideal gyrator for arbitrary spatial and polarization modes.
1 Introduction
@@
Symmetries relate to laws in optics. Noether’s theorem says that a system with a continuous symmetry has a conserved quantity; for example, the frequency (energy) of light is conserved in a medium with the time-translational symmetry. Another kind of symmetries is a discrete one, such as the spatial-inversion and time-reversal symmetry. These symmetries impose restrictions on optical phenomena and play essential roles in optics [1]. A breaking of the space-inversion and/or time-reversal symmetry gives rise to bianisotropies [2, 3], which represent cross-correlated electromagnetic inductions among electric polarizability, magnetization, and electromagnetic fields i.e., magnetoelectric couplings, in constitutive equations. The bianisotropies are classified into the four categories associated with reciprocity and polarization dependence as follows [4]: a reciprocal chiral-type bianisotropy with polarization plane rotation, a reciprocal omega-type without polarization rotation, a non-reciprocal Tellegen-type with polarization rotation, and a non-reciprocal moving-type without polarization rotation. The control of the bianisotroies is a challenge in natural materials with broken symmetries such as multiferroic materials [5, 6, 7]; therefore, man-made structured materials, referred to as metamaterials, are suitable to explore bianisotropies [8, 9, 10, 11, 12]. In particular, metasurfaces with two-dimensional planar structures are good playgrounds to embody a tailor-made bianisotropic medium [13, 14, 15, 16].
@@
In this work, we theoretically and numerically study magnetochiral (MCh) metasurfaces as non-reciprocal moving-type bianisotropic media. The MCh metasurfaces consist of double Z-type gammadion with the broken space-inversion symmetry [17] and perpendicularly magnetized substrate with the broken time-reversal symmetry. The optical MCh effect [18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34] is manifested in a transmission coefficient difference, which is dependent on propagation direction but independent of polarization, due to the moving-type bianisotropy. We calculate transmission and reflection coefficient spectra at normal incidence of microwaves to evaluate effective polarizability tensor [35, 36, 37]. The effective polarizability tensor demonstrates the moving-type bianisotropy as well as the chiral-type bianisotropy and magneto-optical (MO) effect. Furthermore, a composite metasurface consisting of the ordinary and contraposition MCh metasurcfaces realizes pure moving optical media without optical activity and MO effect. The pure moving media show the perfect transmission with a phase difference of $\pi$ depending on propagation direction, bringing about an ideal gyrator [38].
@@
The pure moving optical media are of great interest in terms of development of functional metasurfaces. A similar story can be found in the development of the well-known negative refractive index metamaterials consisting of split-ring resonators (SRRs). A single C-shaped SRR shows the omega-type bianisotropy due to electromagnetic interaction in addition to magnetic permeability [39, 40]. This brings about electromagnetic responses not only to magnetic field but also to electric field. An inverse-C-shaped SRR is thus incorporated in the inner circle of the C-shaped SRR, resulting in vanishment of the omega-type electromagnetic interaction but remaining of the permeability component. This way, the double SRRs construct pure negative magnetic permeability media, bringing about negative refractive index metamaterials. Similarly, the pure moving medium without the optical activity and MO effect obtained in this study is a key to realize an ideal gyrator for arbitrary spatial and polarization modes.
@@
This paper is organized in five sections. Section 2 details theoretical background, including effective polarizability tensor, principle of parameter retrieving, and classification of bianisotropic metasurfaces. Section 3 devotes to numerical calculation setup. Section 4 describes results and discussion of calculated microwave transmission and reflection coefficient spectra of the uniaxial bianisotropic MCh metasurfaces. The extracted effective polarizability tensor of the chiral and MCh metasurfaces is illustrated. We present pure moving optical media by introducing contraposition MCh metasurfaces with the inverse chirality and converse magnetization of the ordinary MCh metasurfaces. The pure moving media result in perfect transmission and phase difference of $\pi$. Section 5 concludes the paper.
2 Theoretical background
2.1 Effective polarizability tensor
@@
The electromagnetic properties of a metasurface, which interacts with normally incident plane waves propagating in the $+z$ direction, can be fully characterized by the following effective polarizability tensor $\hat{\alpha}_{ij}$ defined as
@@
Here we write the plane waves to be, $E,H\propto{\rm e}^{j(-\vec{k}\cdot\vec{r}+\omega t)}$, where $j=\sqrt{-1}$ is the imaginary unit. In this notation, the effective polarizability tensor components can be expressed in a non-dimensional manner to be